The following equations describe an economy (think of C, I, G, etc as being measured in billions and i as a percentage; a 5 percent interest rate implies i = 5) 

C = 0.8 (1 – t) Y

t = 0.25

I = 900 – 50i

G = 800

L = 0.25Y – 62.5.i

M / P = 500


What is the general definition of the IS curve?

What is the equation that describes the IS curve?

What is the equation that describes the LM curve?

What is the general definition of the LM curve?

What are the equilibrium levels of income and the interest rate?




1
Expert's answer
2021-07-01T13:37:42-0400

1) The IS (Investment Savings) curve explains the equilibrium of s product market as it reflects on the relationship between level of income Y and market interest X which appears in the market for goods and services. The IS curve is derived from a simple Keynesian model, however difference is based on total spending and, investment spending relies on interest rate.

b) "Y=C+I+G=0.8(1-t)Y+900-50R+800=0.6Y+900-50R+800=0.8Y+1700.8-50R"

"Y-0.6Y=1700-50R"

"0.4Y=1700-50R"

"Y=4250-125R"

d) LM (liquidity-money) curve describes money market equilibrium that is available when money demand is the same as supply for money. Because money demand is based on rates of interest, there is a money market equilibrium curve.